OpenStudy (anonymous):

A glob of sticky putty is thrown against a slightly open door so that it sticks on impact and gives the door a momentum of 1.4 kg * m/s. A rubber ball with the exactly the same mass and velocity as the sticky putty strikes the door and bounces off back at its initial speed in the opposite direction. What impulse does it impart to the door?

OpenStudy (anonymous):

0. If the ball has the same mass and velocity as the sticky putty then, it hits the door and imparts the same momentum. And you can assume conservation of momentum and say the door imparts that momentum back to the ball as it bounce back off the door so the net momentum change is 0.

OpenStudy (anonymous):

By reversing the ball's momentum, twice that momentum must be transmitted to the door, though conservation of momentum. F = d(mv)/dt force = change in momentum per unit time impulse = F dt = d(mv) = change in momentum.